# m[TODO] #1

### v1

@@ 6.1. Experimental Setting
VSAG maximum_degree $\in\{8,12,16,24,32,36,48,64\}$
pruning_rate $\in\{1.0,1.2,1.4,1.6,1.8,2.0\}$
hnswlib, hnsw(faiss) maximum_degree $\in\{4,8,12,16,24,36,48,64,96\}$
candidate_pool_size $\in\{500\}$
nndescent pruning_prob $\in\{0.0,1.0\}$
leaf_size $\in\{24,36,48\}$
n_neighbors $\in\{10,20,40,60\}$
pruning_degree_multiplier $\in\{0.5,0.75,1.0,1.5,2.0,3.0\}$
@@
scann n_leaves $\in\{100,600,1000,1500,2000\}$
avq_threshold $\in\{0.15,0.2,0.55\}$
dims_per_block $\in\{1,2,3,4\}$
Environment. The experiments are conducted on a server with an Intel(R) Xeon(R) Platinum 8163 CPU @ 2.50GHz and 512GB memory, except that Section 6.2 is evaluated on an AWS r6i.16xlarge machine with hyperthreading disabled. We implement VSAG in C++, and compile it with g++ 10.2.1, -Ofast flag, and AVX-512 instructions enabled. For other baselines, we use the implementations from the official Docker images.
@@ 6.2. Overall Performance
Figure 7 evaluates the recall (Recall@10 and Recall@100) vs. QPS performance of algorithms. We report the best performance of an algorithm under all possible parameter settings. Across all datasets, VSAG can achieve higher QPS with the same recall rate. In addition, VSAG can provide a higher QPS increase on high-dimensional vector datasets, such as GIST1M and OPENAI. In particular, VSAG outperforms hnswlib by 226% in QPS on GIST1M when we ensure $Recall@10=90\%$, it also provides 400% higher QPS than hnswlib on OPENAI when ensuring $Recall@10=80\%$. The reason is that VSAG adopts quantization methods, which can provide significant QPS increase without sacrificing the search accuracy, and they are especially effective for high-dimensional data (Gao and Long, 2024).
6.3. Ablation Study
6.3.1. Cache Miss Analysis
@@ 6.4.1. Tuning Cost
We evaluate the tuning time and memory footprint when varying index-level parameters $m_{c}\in(8,16,24,32)$ and $\alpha_{c}\in(1.0,1.2,1.4,1.6,1.8,2.0)$. The compared method including FIX (fix parameter $m_{c}=32$ and $\alpha_{c}=2.0$, building with hnswlib) and BF (construct indexes of all combination of index-level parameters, building with hnswlib). Table 6 shows results (Mem for memory footprint in gigabytes; Time for building time in hours).
Among the four datasets, the method with minimal construction time and memory usage is FIX, while BF requires the most. This is because baseline methods can only brute-force traverse each parameter configuration and build indexes separately. Additionally, it can dynamically select edges during search using edge labels acquired during construction, enabling parameter adjustment at search time without repeated index construction. This allows VSAG to achieve a 20×20$\times$20 × tuning time saving compared to BF on GIST1M. Moreover, the use of $T$ to cache distances during construction eliminates redundant computations, further accelerating VSAG’s tuning. In terms of memory usage, graphs built with different $m_{c}$ and $a_{c}$ share significant structural overlap. This enables VSAG to compress these graphs via edge labels, resulting in over 20×20$\times$20 × memory footprint reduction during tuning compared to BF. Compared to FIX, VSAG only introduces minimal additional memory costs for storing extra edges and label sets.
@@ 6.5. Evaluation of QLPs Auto-Tuner
Table 6. Comparison of Tuning Cost of ILPs.
Dataset FIX BF VSAG (Ours)
Mem Time Mem Time Mem Time
GIST1M 3.83G 3.20H 89.87G 61.64H 4.07G 2.92H
SIFT1M 0.73G 0.85H 15.48G 30.79H 0.97G 1.86H
GLOVE-100 0.74G 1.22H 15.36G 41.94H 1.03G 2.13H
7. Case Studies and Applications
Large Language Model. In large language model (LLM) applications (Lewis et al., 2020), high-dimensional text embeddings are essential for reducing hallucinations and delivering real-time interactive responses. Modern embedding models, such as text-embedding-3-small [^2] https://platform.openai.com/docs/guides/embeddings, produce 1,536-dimensional vectors, while text-embedding-3-large outputs up to 3,072 dimensions—significantly more than the typical 256-dimensional image embeddings. Effective retrieval over these large embedding vectors demands both precise distance computation and careful memory management. VSAG ’s dedicated distance computation and memory access optimizations excel at handling these high-dimensional workloads.
@@ 8. Related Work
The mainstream methods for vector retrieval can be divided into two categories: space partitioning-based and graph-based. Graph-based algorithms (Malkov and Yashunin, 2020a; Jayaram Subramanya et al., 2019a; Fu et al., 2019a; Fu et al., 2022; Lu et al., 2021; Peng et al., 2023a; Azizi et al., 2023; Li et al., 2020) can ensure high recall with practical efficiency, e.g., HNSW (Malkov and Yashunin, 2020a), NSG (Fu et al., 2019b), VAMANA (Jayaram Subramanya et al., 2019a), and $\tau$-MNG (Peng et al., 2023b). These methods build a proximity graph where each node is a base vector and edges connect pairs of nearby vectors. During a vector search, they greedily move towards the query vector to identify its nearest neighbors. Our VSAG framework can adapt to graph-based algorithms mentioned above, to improve the performance in production.
@@ 9. Conclusion
In this work, we present VSAG, an open-source framework for ANNS that can be applied to most of the graph-based indices. To address the challenge of high memory access costs, we employ software-based prefetch, deterministic access greedy search, and PRS to significantly reduce the cache miss rate. For the challenge of high parameter tuning costs, we propose a three-level parameter tuning mechanism that automatically adjusts different parameters based on their tuning complexity. To tackle the issue of high distance computation costs, we combine quantization and selective re-ranking to integrate low- and high-precision distance computations, thereby enhancing search efficiency without sacrificing effectiveness. Experiments on real-world datasets demonstrate that VSAG outperforms baselines in both retrieval performance and parameter tuning cost.
@@ Appendix A Detail Explanation of Deterministic Access Greedy Search
In VSAG, the insertion process for each new point $x_{i}$ (Lines 3–4) can be divided into three steps. First, a search is conducted on the graph to obtain the approximate nearest neighbors $\mathrm{ANN}_{k}(x_{i})$ and their corresponding distances $T_{i}$ (Line 5). Next, pruning is performed on $\mathrm{ANN}_{k}(x_{i})$, and the edges along with their labels are added to the graph index (Line 6). The first step, the search process, is similar to that of other graph-based index constructions. However, the second step, the pruning process, differs from other indices. Instead of performing actual pruning, VSAG assigns labels. Specifically, VSAG assigns a label $L_{i,j}$ to each edge $G_{i,j}$ in graph (see Algorithm 2). Note that the label is used to indicate whether an edge exists under a certain construction parameter. Given $\alpha_{s}$, if $L_{i,j}\leq\alpha_{s}$, it indicates that the edge $G_{i,j}$ exists in the graph constructed using $\alpha_{s}$ or larger value as the pruning rate. During the search and pruning processes, only edges satisfying $L_{i,j}\leq\alpha_{s}$ need to be considered to retrieve the graph corresponding to the given construction parameters, which is proven in Theorem 4.3.

### v3

@@ 6.1. Experimental Setting
VSAG maximum_degree $\in\{8,12,16,24,32,36,48,\textbf{64}\}$
pruning_rate $\in\{\textbf{1.0},1.2,1.4,1.6,1.8,2.0\}$
hnswlib, hnsw(faiss) maximum_degree $\in\{4,8,12,16,24,36,48,\textbf{64},96\}$
candidate_pool_size $\in\{\textbf{500}\}$
nndescent pruning_prob $\in\{\textbf{0.0},1.0\}$
leaf_size $\in\{24,\textbf{36},48\}$
n_neighbors $\in\{10,20,\textbf{40},60\}$
pruning_degree_multiplier $\in\{0.5,0.75,1.0,1.5,\textbf{2.0},3.0\}$
@@
scann n_leaves $\in\{100,600,1000,1500,\textbf{2000}\}$
avq_threshold $\in\{0.15,0.2,\textbf{0.55}\}$
dims_per_block $\in\{1,2,3,\textbf{4}\}$
Environment. The experiments are conducted on a server with an Intel(R) Xeon(R) Platinum 8163 CPU @ 2.50GHz and 512GB memory, except that Section 6.2 is evaluated on an AWS r6i.16xlarge machine with hyperthreading disabled, and Section 6.3 is evaluated on a server with 4 AMD EPYC 7T83 64-Core Processors and 2TB memory. We implement VSAG in C++, and compile it with g++ 10.2.1, -Ofast flag, and AVX-512 instructions enabled. For baselines, we use the implementations from the official Docker images.
@@ 6.2. Overall Performance
Figure 6 evaluates the recall (Recall@10 and Recall@100) vs. QPS performance of algorithms. We report the best performance of an algorithm under all possible parameter settings. Across all datasets, VSAG can achieve higher QPS with the same recall rate. In addition, VSAG can provide a higher QPS increase on high-dimensional vector datasets, such as GIST1M and OPENAI. In particular, VSAG outperforms hnswlib by 226% in QPS on GIST1M when fixing $Recall@10=90\%$, it also provides $\sim$400higher QPS than hnswlib on OPENAI when fixing $Recall@10=80\%$. The reason is that VSAG adopts quantization methods, which can provide significant QPS increase without sacrificing the search accuracy, and they are especially effective for high-dimensional data (Gao and Long, 2024).
@@
6.3. Scalability Performance
(a) ANT-INTERNET ($\sim$10M)
(b) MSMARCO ($\sim$100M)
Figure 7. Performance Comparison on Large-Scale Datasets
As shown in Figure 7, we present the performance comparison of VSAG with four baselines is on two large-scale datasets: ANT-INTERNET($\sim$10M) and MSMARCO($\sim$100M). Here, ANT-INTERNET is an internal text dataset from Ant Group. We made our best effort to construct the well-tuned indexes for the MSMARCO dataset using 256 threads. The performance of hnsw (faiss) is similar to that of hnswlib, while the performance of faiss-ivf is significantly worse than that of faiss-ivfpqfs. Additionally, nndescent failed to complete index construction within two days on the MSMARCO dataset. As a result, we do not include these methods in the results.
For faiss-ivfpqfs, we followed the META-FAISS guidebook to tune its parameters. To strike a balance between construction time and performance, we use 1 million centroids and employed PQ with $256\times 4$ bits. For the other algorithms, we use the parameters highlighted in bold in Table 4.
For VSAG, the construction process took 15.37 hours with a memory footprint of 463GB on the MSMARCO dataset. As the results shown, at Recall@10=99% VSAG achieves a significant improvement in QPS compared to hnswlib (increased from 180 to 467), which is a 2.59×2.59$\times$2.59 × performance boost. Similarly, on the ANT-INTERNET dataset, at Recall@10=96% VSAG demonstrates a remarkable enhancement in QPS rising from 659 to 1421, which is 2.15×2.15$\times$2.15 × QPS of hnswlib.
6.4. Ablation Study
6.4.1. Cache Miss Analysis
@@ 6.5.1. Tuning Cost
Table 6. Time Cost Breakdown in Each Phase.
Dataset Algorithm Offline Construction and Tuning Phase Online Search Phase (1,000 queries)
Build Index ILP Tuning ELP Tuning QLP Training QLP Tuning Search with Query
SIFT1M VSAG 5998.564s 68.185s 10.293s 50.768s $\sim$0.001s 0.217s
hnswlib 30.79 hours (18×18$\times$18 ×) - 0.370s
GIST1M VSAG 11085.626s 119.813s 25.397s 149.467s $\sim$0.001s 1.587s
hnswlib 61.64 hours (19×19$\times$19 ×) - 4.807s
As shown in Table 6, we present the time consumption of VSAG across different phases on the SIFT1M and GIST1M datasets, varying index-level parameters $m_{c}\in(8,16,24,32)$ and $\alpha_{c}\in(1.0,1.2,1.4,1.6,1.8,2.0)$. Here, we report the total time for 1,000 queries at Recall@10 = 99%. During the offline phase, the tuning of ILP (i.e., 68s) and ELP (i.e., 10s) takes the longest time, as it involves adjusting a large number of parameters and performing actual searches to select the optimal parameters. However, the total time spent on all tuning (i.e., 128s) processes is still significantly lower than the time required for building index (i.e., 5998s). In contrast, hnswlib requires over 30 hours to repeatedly construct indexes with different ILP during its tuning process. In the online phase, the tuning of QLP introduces only 0.001s overhead because the decision tree only relies on a small number of features. This overhead is almost negligible compared to the search cost (i.e., 0.217s). Even when including the tuning overhead of QLP, VSAG’s overall search cost remains lower than that of hnswlib (i.e., 0.37s).
@@ 6.6. Evaluation of QLPs Auto-Tuner
7. Case Study
In Ant Group, vector retrieval capabilities are provided through a distributed vector database system, which is designed to support real-time queries with highly availability and high scalability. For storage, the vector database system splits the billion-level dataset into subsets managed by the LSM-Tree structure (Guo et al., 2022), and each subset is referred to as a segment. Our VSAG library focuses solely on indexing and retrieval performance. Therefore, when VSAG is integrated into a vector database, we separately construct a VSAG index for each segment. Then, for a coming query, it is executed on indexes on Segments in parallel, and the vector database system merges the results from segments as the final result.
For the deployment of the vector database system, we use query nodes to manage the segments. Here, a query node refers to a server that responds to online requests. To minimize the impact of query node failures and avoid prolonged recovery times, the size and number of segments maintained by a single distributed node are limited. As data volume grows, scalability is achieved by adding more nodes horizontally. This approach is also widely adopted in vector databases such as Milvus (Guo et al., 2022) and Vald (Vald, 2021).
For example, in an image search scenario involving approximately 10 billion images in Ant Group, each image is embedded into a 512-dimensional vector and stored in a distributed vector database cluster. The cluster is configured such that each segment contains approximately 10 million rows of data. Each query node can host up to 4 segments, and each query node is deployed on an machine instance with a 16-core CPU and 80GB memory. The entire cluster consists of approximately 400 such machine instances. By using the VSAG index, the average latency in this scenario is reduced from 3.0ms to 1.1ms per segment, and the upper limit of QPS throughput is increased by 2.65×2.65$\times$2.65 ×, compared to using the open-source hnswlib algorithm.
@@ 8. Related Work
The mainstream methods for vector retrieval can be divided into two categories: space partitioning-based and graph-based. Graph-based algorithms (Malkov and Yashunin, 2020a; Jayaram Subramanya et al., 2019a; Fu et al., 2019a; Fu et al., 2022; Lu et al., 2021; Peng et al., 2023a; Azizi et al., 2023; Li et al., 2020; Yang et al., 2025) can ensure high recall with practical efficiency, e.g., HNSW (Malkov and Yashunin, 2020a), NSG (Fu et al., 2019b), VAMANA (Jayaram Subramanya et al., 2019a), and $\tau$-MNG (Peng et al., 2023b). These methods build a proximity graph where each node is a base vector and edges connect pairs of nearby vectors. During a vector search, they greedily move towards the query vector to identify its nearest neighbors. Our VSAG framework can adapt to graph-based algorithms mentioned above, to improve the performance in production.
@@ 9. Conclusion
In this work, we present VSAG, an open-source framework for ANNS that can be applied to most of the graph-based indexes. VSAG employs software-based prefetch, deterministic access greedy search, and PRS to significantly reduce the cache miss rate. VSAG has a three-level parameter tuning mechanism that automatically adjusts different parameters based on their tuning complexity. VSAG combines quantization and selective re-ranking to integrate low- and high-precision distance computations. Experiments on real-world datasets demonstrate that VSAG outperforms baselines.
10. acknowledgment
This work was supported by Ant Group. Heng Tao Shen and Peng Cheng are partially supported by the Fundamental Research Funds for the Central Universities. Xuemin Lin is supported by NSFC U2241211.
@@ Appendix A Detail Explanation of Deterministic Access Greedy Search
In VSAG, the insertion process for each new point $x_{i}$ (Lines 3–4) can be divided into three steps. First, a search is conducted on the graph to obtain the approximate nearest neighbors $\mathrm{ANN}_{k}(x_{i})$ and their corresponding distances $T_{i}$ (Line 5). Next, pruning is performed on $\mathrm{ANN}_{k}(x_{i})$, and the edges along with their labels are added to the graph index (Line 6). The first step, the search process, is similar to that of other graph-based index constructions. However, the second step, the pruning process, differs from other indexes. Instead of performing actual pruning, VSAG assigns labels. Specifically, VSAG assigns a label $L_{i,j}$ to each edge $G_{i,j}$ in graph (see Algorithm 2). Note that the label is used to indicate whether an edge exists under a certain construction parameter. Given $\alpha_{s}$, if $L_{i,j}\leq\alpha_{s}$, it indicates that the edge $G_{i,j}$ exists in the graph constructed using $\alpha_{s}$ or larger value as the pruning rate. During the search and pruning processes, only edges satisfying $L_{i,j}\leq\alpha_{s}$ need to be considered to retrieve the graph corresponding to the given construction parameters, which is proven in Theorem 4.2.
@@ Appendix C Proof of Theorem
Appendix E Tuning Cost of ILPs
Table 8. Comparison of Tuning Cost of ILPs.
Dataset FIX BF VSAG (Ours)
Mem Time Mem Time Mem Time
GIST1M 3.83G 3.20H 89.87G 61.64H 4.07G 2.92H
SIFT1M 0.73G 0.85H 15.48G 30.79H 0.97G 1.86H
GLOVE-100 0.74G 1.22H 15.36G 41.94H 1.03G 2.13H
We evaluate the tuning time and memory footprint when varying index-level parameters $m_{c}\in(8,16,24,32)$ and $\alpha_{c}\in(1.0,1.2,1.4,$ 1.6,1.8,2.0)1.6,1.8,2.0)1.6 , 1.8 , 2.0 ). The compared method including FIX (fix parameter $m_{c}=32$ and $\alpha_{c}=2.0$, building with hnswlib) and BF (construct indexes of all combination of index-level parameters, building with hnswlib). Table 8 shows results (Mem for memory footprint in gigabytes; Time for building time in hours).
Among the four datasets, the method with minimal construction time and memory usage is FIX, while BF requires the most. This is because baseline methods can only brute-force traverse each parameter configuration and build indexes separately. Additionally, it can dynamically select edges during search using edge labels acquired during construction, enabling parameter adjustment at search time without repeated index construction. This allows VSAG to achieve a 19×19$\times$19 × tuning time saving compared to BF on GIST1M. Moreover, the use of $T$ to cache distances during construction eliminates redundant computations, further accelerating VSAG’s tuning. In terms of memory usage, graphs built with different $m_{c}$ and $a_{c}$ share significant structural overlap. This enables VSAG to compress these graphs via edge labels, resulting in over 19×19$\times$19 × memory footprint reduction during tuning compared to BF. Compared to FIX, VSAG only introduces minimal additional memory costs for storing extra edges and label sets.
Appendix F Applications
Large Language Model. In large language model (LLM) applications (Lewis et al., 2020), high-dimensional text embeddings are essential for reducing hallucinations and delivering real-time interactive responses. Modern embedding models, such as text-embedding-3-small, produce 1,536-dimensional vectors, while text-embedding-3-large outputs up to 3,072 dimensions—significantly more than the typical 256-dimensional image embeddings. Effective retrieval over these large embedding vectors demands both precise distance computation and careful memory management. VSAG ’s dedicated distance computation and memory access optimizations excel at handling these high-dimensional workloads.
Figure 10. Truncated scalar quantization.
Appendix G Improving Quantization Precision
Scalar quantization (SQ) compresses floating-point vectors by mapping 32-bit float values to lower-bit integer representations in each dimension. In SQ-b (where b denotes bit-width), the dynamic range is uniformly partitioned into $2^{b}$ intervals. For SQ4 (4-bit) on vector $x$, this creates 16 intervals over $[\min(x),\max(x)]$. Values of each dimension of $x$ within each interval are encoded as their corresponding integer indices. However, practical implementations face critical range estimation challenges: direct use of observed min/max values proves suboptimal when data outliers existing.
Example 0.